### v1

# Prefix Sums via Kronecker Products

Abstract
In this work, we revisit prefix sums through the lens of linear algebra. We describe an identity that decomposes triangular all-ones matrices as a sum of two Kronecker products, and apply it to design recursive prefix sum algorithms and circuits. Notably, the proposed family of circuits is the first one that achieves the following three properties simultaneously: (i) zero-deficiency, (ii) constant fan-out per-level, and (iii) depth that is asymptotically strictly smaller than $2\log(n)$ for input length $n$. As an application, we show how to use these circuits to design quantum adders with $1.893\log(n)+O(1)$ Toffoli depth, $O(n)$ Toffoli gates, and $O(n)$ additional qubits, improving the Toffoli depth and/or Toffoli size of existing constructions.
I Introduction
Given a sequence of $n$ elements $\bm{x}(0),\bm{x}(1),\ldots,\bm{x}(n-1)$ and an associative binary operator $\circ$, the “prefix problem”, also known as “prefix sum” or “scan”, is defined as the sequence:
$\bm{y}(i)=\bm{x}(0)\circ\bm{x}(1)\circ\ldots\circ\bm{x}(i),\quad 0\leq i\leq n-1.$ (1)
Prefix sum is a fundamental concept in algorithm analysis with numerous applications. It plays a central role in the design of arithmetic processors, such as binary adders [1, 2] and multipliers [3], even on quantum computers [4, 5]. It is widely used as a building block for parallel graph algorithms [6], matrix arithmetics [7], and, more recently, language models [8]. Due to the wide applicability, prefix sum algorithms have been studied for decades, and their theoretical properties are well-understood.
The prevailing model to analyze prefix sums is the circuit model; see e.g. [3]. For a sequence of length $n$, a prefix circuit is a directed acyclic graph where the nodes indicate binary operations $\circ$. The size $S(n)$ is the total number of nodes, the depth $D(n)$ is the length of the longest path (number of edges) from an input to an output, and the fan-out is the maximum number of outgoing edges of any node. Simplicity of the circuit constructions is also desirable for efficient implementations. It is typically quantified mathematically with the so-called uniformity of the circuit family, which measures how much time or space is required by an algorithm to either construct the circuit or to verify its properties.
A plethora of sophisticated prefix circuit constructions have been proposed in the literature with logarithmic depth [1, 9, 2, 10, 11, 12, 13, 14, 15, 16, 17] (see Table 1). Sklansky [1] described a circuit that achieves the optimal depth $D(n)=\log(n)$, but the size increases to $n\log(n)/2$. Ladner and Fischer [9] proposed a recursive construction that can reduce the size by slightly increasing the depth. Brent and Kung [2] described a recursive circuit in the context of parallel adders. Snir [10] proved a tight optimality condition for the sum of depth and size of any prefix circuit, specifically, the following inequality:
$D(n)+S(n)\geq 2n-2.$ (2)
Circuit families that satisfy this bound with equality are said to have zero-deficiency, and they have attracted significant attention since Snir’s discovery. The so-called “LYD” family of circuits [11] improved the depth of Snir’s construction while maintaining zero-deficiency, and Lin and Shih further improved the minimum achievable depth [12]. Lin and co-authors published a series of further developments, focusing on zero-deficiency parallel prefix circuits with bounded fan-out (at most $4$) [13, 14, 15, 16, 17]. In a landmark work, Zhu, Cheng, and Graham [18] provided a simplified proof of Snir’s bound, and they also proved that there exist no zero-deficiency circuits with depth less than:
$D_{\min}(n):=\min\{t:F(t)\geq n+1\}-3,$ (3)
where $F(t)$ is the $t$-th Fibonacci number. In addition, they described a circuit family that achieves this lower bound for every $n$ (see also Table 1).
While the focus of this work is on circuits with bounded fan-in, here we also mention several works that have focused in unbounded fan-in circuits. A seminal result in this literature is that if the underlying semi-group is group free then there there exist constant depth and almost linear size circuits  [19, 3, 20]. For example, in [3] it is shown that, for the input length $n$, there exist prefix circuits of depth $4$ and size $2n\log^{*}(n)$.
Table 1: Prefix circuits.
Size $S(n)$ Depth $D(n)$ Fan-out Deficiency Reference
n1n-1 n1n-1 22 0 Serial circuit.(2)
$\tfrac{n}{2}\log(n)$ $\log(n)$ $n/2$ $O(n\log(n))$ Sklansky [1].(1)(2)
$n\lceil\log(n)\rceil-n+1$ $\log(n)+O(1)$ 22 $O(n\log(n))$ Kogge–Stone [21],[3, Sec. 3.1, Fig. 1].(2)
2n22n-2 $\log(n)+O(1) 22$ $O(\log(n))$ Cyclic Reduction [3, Sec. 3.1, Fig. 5].(2)
$2(1+\tfrac{1}{2^{k}})n-o(n)-k$ $\log(n)+k$ $\lfloor\tfrac{n+2^{k}-1}{2^{k+1}}\rfloor+k$ $O(n)$ Ladner–Fischer [9], $0\leq k\leq\log(n)$.(1)
$2n-\log(n)-2$ $\log(n)-1 22$ $O(\log(n))$ Brent–Kung [2].(1)(2)
2n2D(n)2n-2-D(n) $2\log(n)+O(1)$ $\lceil\log(n)\rceil+1$ 0 Snir [10].
2n2D(n)2n-2-D(n) $2\log(n)+O(1)$ $2\lceil\log(n)\rceil-2$ 0 LYD [11], $n\geq 9$.
2n2D(n)2n-2-D(n) $2\log(n)+O(1)$ $\lceil\log(n)\rceil+1$ 0 SL [12], $n\geq 12$.
2n2D(n)2n-2-D(n) $2\log(n)+O(1)$ 44 0 H4 [13], Z4 [14], WE4 [15], SU4 [17].
2n2D(n)2n-2-D(n) $\mathbf{1.4401}\log(n)+O(1)$ $D(n)$ 0 ZCG [18].
2n2D(n)2n-2-D(n) $2\log(n)+O(1)$ $\mathbf{3}$ 0 This work: Thm. 2 for s=2s=2.(2)
2n2D(n)2n-2-D(n) $\mathbf{1.8928}\log(n)+O(1)$ 44 0 This work: Thm. 2 for s=3s=3.(2)
(1) The size and depth is reported only when $n$ is a power of two.
(2) The circuit construction is LOGTIME-uniform.
Contributions.
In this work, we propose a new approach of constructing prefix sum circuits and algorithms by viewing the problem from a linear algebraic perspective. Specifically, recall that the prefix sum of a vector $\bm{x}$ can be written as $\bm{y}=\bm{L}\bm{x}$, where $\bm{L}$ is the lower triangular all-ones matrix. As we note in Theorem 1, $\bm{L}$ admits a two-term Kronecker product decomposition (see also Figure 1). This allows us to describe many different types of prefix algorithms, both existing and new ones, by exploiting the properties of Kronecker products.
Figure 1: Illustration of Kronecker decomposition (Theorem 1).
By carefully organizing the underlying operations, we obtain a new family of recursive prefix circuits, whose properties are summarized in Table 1 (see also Theorem 2). These circuits are parametrized by an integer $s\in[2,n/2]$ which corresponds to the “block-size” of the recursion. Different choices for $s$ yield circuits with different properties. We highlight the following cases:
  • For $s=2$, the proposed circuit has the lowest known fan-out compared to all known zero-deficiency constructions with depth $2\log(n)+O(1)$.
  • For $s=3$, it has the lowest depth $(\approx 1.893\log(n)+O(1))$ between all circuits with constant fan-out.
Importantly, the proposed circuits are quite simple to describe. Specifically, they satisfy the so-called LOGTIME-uniformity, which means that the connections between the nodes of a circuit can be verified in $O(\log(n))$ time for a sequence of length $n$ [22]. To the best of our knowledge this is the first LOGTIME-uniform prefix circuit family that achieves depth smaller than $2\log(n)+O(1)$.
As a direct application, we show how to use the proposed circuits in the context of quantum adders. Evidently, the proposed quantum adder achieves lower Toffoli depth and/or size compared to existing adders, as summarized in Table 3.
Notation.
Matrices and vectors are denoted with capital and small Latin letters, respectively, in bold font. All vectors are considered column vectors and $\bm{1}_{n}$ is the all-ones vector with length $n$. We denote by $\bm{U}_{s}$ (and $\bm{L}_{s}$) the upper (and lower) triangular all-ones matrix of size $s\times s$. $\bm{I}_{n}$ is the identity matrix of size $n\times n$, and $\bm{e}_{i}$ is its $i$-th column. $\bm{U}^{-}_{n}:=\bm{U}_{n}-\bm{I}_{n}$, and $\bm{L}^{-}_{n}:=\bm{L}_{n}-\bm{I}_{n}$ are the strictly upper- and lower-triangular all-ones matrices, respectively. For non-negative integers $i,j,k$ we denote $[i:j]$ the set $\{i,i+1,\ldots,j\}$ and $[i:k:j]$ the set $\{i,i+k,i+2k,\ldots,i+mk\}$, where $m$ is the largest integer such that $i+mk\leq j$. We use zero-based indexing for matrices and vectors. $\bm{A}(i,j)$ is the element of $\bm{A}$ in row $i$, column $j$. For a vector $\bm{x}$, $\bm{x}(i)$ is the $(i+1)$-th entry, and $\bm{x}(i:j)$ is the subvector $(\bm{x}(i),\bm{x}(i+1),\dots,\bm{x}(j))$. For $1<s<n$, the operator $\operatorname{mat}_{s}:\mathbb{R}^{n}\to\mathbb{R}^{s\times\lceil n/s\rceil}$ returns a matrix: $\operatorname{mat}_{s}(\bm{x})=\begin{pmatrix}\bm{x}_{0}&\bm{x}_{1}&\ldots&\bm{x}_{\lceil\frac{n}{s}\rceil}\end{pmatrix}$, where $\bm{x}$ is padded with zeros if $s$ does not divide $n$ and $\bm{x}_{i}:=\bm{x}(is:is+s-1)$. $\operatorname{vec}(\bm{X})$ stacks the columns of $\bm{X}$ in a single column vector. We denote $h_{s}(n):=\lceil\frac{n}{s}\rceil-1$, $h_{s}^{(k)}(n)=\underbrace{h_{s}(h_{s}(\ldots h_{s}}_{k\text{ times}}(n)\ldots))$, $h_{s}^{*}(n)=\arg\max_{j}\left\{h_{s}^{(j)}(n)>s\right\}$, and $r_{s}(n):=h^{(h^{*}_{s}(n))}_{s}(n)\in[s]$ is the “remainder”.
II Kronecker products
In this section, we prove that triangular all-ones matrices can be written as the sum of two Kronecker products of triangular all-ones matrices. We first recall some useful properties of Kronecker products.
Let $\bm{A}\in\mathbb{R}^{m\times n}$ and $\bm{B}\in\mathbb{R}^{p\times q}$ be two matrices, and denote $a_{i,j}:=\bm{A}(i,j)$. The Kronecker product of $\bm{A}$ and $\bm{B}$, denoted by $\bm{A}\otimes\bm{B}$, is the $mp\times nq$ block matrix:
$\bm{A}\otimes\bm{B}$ $=\begin{bmatrix}a_{0,0}\bm{B}&a_{0,1}\bm{B}&\cdots&a_{0,n-1}\bm{B}\\ a_{10}\bm{B}&a_{11}\bm{B}&\cdots&a_{1,n-1}\bm{B}\\ \vdots&\vdots&\ddots&\vdots\\ a_{m-1,0}\bm{B}&a_{m-1,1}\bm{B}&\cdots&a_{m-1,n-1}\bm{B}\end{bmatrix}$.
We frequently use the following two Kronecker properties, see [23, Equations $2.4$ and $2.10$]:
Transpose
$(\bm{A}\otimes\bm{B})^{\top}=\bm{A}^{\top}\otimes\bm{B}^{\top}$,
Mixed product
$\operatorname{vec}(\bm{A}\bm{X}\bm{B})=(\bm{B}^{\top}\otimes\bm{A})\operatorname{vec}(\bm{X})$.
Theorem 1.
Fix three integers $n,n_{1},n_{2}>1$, such that $n=n_{1}n_{2}$. The lower triangular all-ones matrix $\bm{L}_{n}$ can be decomposed as follows:
$\bm{L}_{n}$ $=\bm{L}_{n_{1}}\otimes\bm{L}_{n_{2}}+\bm{L}^{-}_{n_{1}}\otimes\bm{U}^{-}_{n_{2}}$
$=\bm{I}_{n_{1}}\otimes\bm{L}_{n_{2}}+\bm{L}^{-}_{n_{1}}\otimes\bm{1}_{n_{2}}\bm{1}^{\top}_{n_{2}}.$ (4)
Equivalently, $\bm{U}_{n}=\bm{U}_{n_{1}}\otimes\bm{U}_{n_{2}}+\bm{U}^{-}_{n_{1}}\otimes\bm{L}^{-}_{n_{2}}$.
Proof.
To prove Equation (1), it suffices to show that the following equality holds:
$\bm{U}_{n}=\bm{U}_{n_{1}}\otimes\bm{1}_{n_{2}}\bm{1}^{\top}_{n_{2}}-\bm{I}_{n_{1}}\otimes\bm{L}_{n_{2}}^{-}.$ (5)
Indeed,
$\bm{U}_{n_{1}}$ $\otimes\bm{1}_{n_{2}}\bm{1}^{\top}_{n_{2}}-\bm{I}_{n_{1}}\otimes\bm{L}_{n_{2}}^{-}$
$=\bm{U}_{n_{1}}\otimes\left(\bm{U}_{n_{2}}+\bm{L}^{-}_{n_{2}}\right)-\bm{I}_{n_{1}}\otimes\bm{L}_{n_{2}}^{-}$
$=\bm{U}_{n_{1}}\otimes\bm{U}_{n_{2}}+\left(\bm{U}_{n_{1}}-\bm{I}_{n_{1}}\right)\otimes\bm{L}_{n_{2}}^{-}$
$=\bm{U}_{n_{1}}\otimes\bm{U}_{n_{2}}+\bm{U}^{-}_{n_{1}}\otimes\bm{L}_{n_{2}}^{-},$
where, in the first equality, we decomposed the all-ones matrix into its upper triangular and strictly lower triangular all-ones parts, and in the next two equalities, we rearranged terms.
Next, we prove that the right-hand side (RHS) of Equation (5) equals $\bm{U}_{n}$ by first showing that its upper triangular part is all-ones, and then, showing that its strictly lower triangular part is all-zeros.
First, notice that the matrix $\bm{I}_{n_{1}}\otimes\bm{L}_{n_{2}}^{-}$ is strictly lower triangular by construction. Moreover, all entries in the upper triangular part of the matrix $\bm{U}_{n_{1}}\otimes\bm{1}_{n_{2}}\bm{1}^{\top}_{n_{2}}$ is one. These two statements imply that the entries of the upper triangular part of the RHS of Equation (5) are equal to one.
Second, it suffices to show that the strictly lower triangular part of the matrix in the RHS of Equation (5) is zero. By construction, all strictly lower triangular blocks of size $n_{2}\times n_{2}$ of the RHS of Equation (5) are zero matrices, so we only need to show that all the $n_{2}\times n_{2}$ block matrices on the main diagonal of the RHS equal $\bm{U}_{n_{2}}$. Consider any diagonal block of size $n_{2}\times n_{2}$ of the RHS, and notice that both Kronecker matrices share the same sizes $n_{1}$ and $n_{2}$ in their product. Hence, the first matrix summand is $\bm{1}_{n_{2}}\bm{1}^{\top}_{n_{2}}$ and the second matrix summand is $\bm{L}_{n_{2}}^{-}$ which gives $\bm{1}_{n_{2}}\bm{1}^{\top}_{n_{2}}-\bm{L}_{n_{2}}^{-}=\bm{U}_{n_{2}}$. ∎
These decompositions can also be extended to strictly upper or lower triangular all-ones matrices by subtracting the identity matrix on both sides of Equation (1).
Given Theorem 1, the following linear algebraic formulation of prefix sum follows. Let $\bm{x}$ be a vector of size $n$ with $n=n_{1}n_{2}$ as in Theorem 1, and let $\bm{X}=\operatorname{mat}_{n_{2}}(\bm{x})\in{\mathbb{R}^{n_{2}\times n_{1}}}$. It holds that:
$\bm{L}_{n}\bm{x}$ $=\left(\bm{I}_{n_{1}}\otimes\bm{L}_{n_{2}}+\bm{L}^{-}_{n_{1}}\otimes\bm{1}_{n_{2}}\bm{1}^{\top}_{n_{2}}\right)\bm{x}$
$=\left(\bm{I}_{n_{1}}\otimes\bm{L}_{n_{2}}\right)\operatorname{vec}(\bm{X})+\left(\bm{L}^{-}_{n_{1}}\otimes\bm{1}_{n_{2}}\bm{1}^{\top}_{n_{2}}\right)\operatorname{vec}(\bm{X})$
$=\operatorname{vec}\left(\bm{L}_{n_{2}}\bm{X}\bm{I}_{n_{1}}\right)+\operatorname{vec}\left(\bm{1}_{n_{2}}\bm{1}_{n_{2}}^{\top}\bm{X}\bm{L}^{-\top}_{n_{1}}\right)$
$=\operatorname{vec}\left(\bm{L}_{n_{2}}\bm{X}\right)+\operatorname{vec}\left(\bm{1}_{n_{2}}\bm{1}_{n_{2}}^{\top}\bm{X}\bm{U}^{-}_{n_{1}}\right),$ (6)
where we used Theorem 1, linearity, and the Kronecker mixed-product property. Interestingly, this formulation lends itself for recursive evaluations. The first term $(\bm{L}_{n_{2}}\bm{X})$ can be evaluated as $n_{1}$ prefix sums of length $n_{2}$. The second term, $\bm{1}_{n_{2}}(\bm{1}_{n_{2}}^{\top}\bm{X})\bm{U}^{-}_{n_{1}}$, is an exclusive prefix sum of length $n_{1}$. By tuning the parameters $n_{1}$ and $n_{2}$, and by choosing evaluation strategies appropriately, we can recover known algorithms and circuits (see Table 2), as well as to design new ones, as discussed next.
Table 2: Three examples of prefix sum algorithms that are derived by (recursively) evaluating a Kronecker product expression of $\bm{L}_{n}$ or $\bm{U}_{n}$.
Expression of $\bm{L}_{n}$ Recursive step(s) Ref.
$\bm{I}_{\frac{n}{2}}\otimes\bm{L}_{2}+\bm{L}^{-}_{\frac{n}{2}}\otimes\bm{1}_{2}\bm{1}^{\top}_{2}$ $\left(\bm{1}_{2}^{\top}\operatorname{mat}_{2}(\bm{x})\right)\bm{U}_{\frac{n}{2}}$ [2]
$\bm{I}_{2}\otimes\bm{L}_{\frac{n}{2}}+\bm{L}_{2}^{-}\otimes\bm{1}_{\frac{n}{2}}\bm{1}^{\top}_{\frac{n}{2}}$ $\bm{L}_{\frac{n}{2}}\bm{x}_{0}$ and $\left(\bm{1}_{2}^{\top}\operatorname{mat}_{2}(\bm{x}_{0})\right)\bm{U}_{\frac{n}{4}}$ [9]
where $\bm{x}_{0}:=\operatorname{mat}_{\frac{n}{2}}(\bm{x})\bm{e}_{0}$
$\bm{I}_{\frac{n}{s}}\otimes\bm{L}_{s}+\bm{L}^{-}_{\frac{n}{s}}\otimes\bm{1}_{s}\bm{1}^{\top}_{s}$ $\left(\bm{1}_{s}^{\top}\operatorname{mat}_{s}(\bm{x})\right)\bm{U}_{\frac{n}{s}}$ [24]
$\bm{I}_{\frac{n}{s}}\otimes\bm{L}_{s}+\bm{L}^{-}_{\frac{n}{s}}\otimes\bm{1}_{s}\bm{1}^{\top}_{s}$ $\left(\bm{1}_{s}^{\top}\operatorname{mat}_{s}(\bm{x})\right)\begin{pmatrix}\bm{U}_{\lceil\frac{n}{s}\rceil-1}&0\\ 0&0\end{pmatrix}$ Thm. 2
III Constructing prefix circuits
We start by showing that the Brent–Kung algorithm [25] can be described as special cases of Eq. (1). Here we set $n_{2}=2$ and $n_{1}=n/2$ (assuming that $n$ is even). Brent–Kung executes the following steps.
  1. 1.
    First, it computes partial prefix sums of size $2$, i.e. $\bm{L}_{2}\bm{X}$, where $\bm{X}=\operatorname{mat}_{2}(\bm{x})\in\mathbb{R}^{2\times\frac{n}{2}}$. In linear algebra notation we write: $\begin{pmatrix}1&0\\ 1&1\end{pmatrix}\bm{X}=\begin{pmatrix}\bm{X}(0,:)\\ \bm{1}_{2}^{\top}\bm{X}\end{pmatrix}$.
  2. 2.
    Next, the prefix sum of $\bm{w}=\bm{1}_{2}^{\top}\bm{X}$ is computed recursively. This can be written as $(\bm{1}_{2}^{\top}\bm{X})\bm{U}_{\frac{n}{2}}$.
  3. 3.
    Finally, the elements of $\bm{w}=\bm{1}_{2}^{\top}\bm{X}$ are used to complete the elements of the first row of $\bm{X}$ (except the first element). This can be written as:
    $\bm{X}(0,1:\tfrac{n}{2}-1)+\bm{w}^{\top}(0:\tfrac{n}{2}-2).$ (7)
Ultimately, the entire algorithm returns $\operatorname{vec}(\bm{Y})$ where:
$\bm{Y}=\begin{pmatrix}\bm{X}(0,:)\\ \underbrace{\underbrace{\bm{1}_{2}^{\top}\bm{X}}_{\text{Step 1}}\bm{U}_{\frac{n}{2}}}_{\text{Step 2}}\end{pmatrix}\underbrace{+\begin{pmatrix}0&\left(\bm{1}_{2}^{\top}\bm{X}\bm{U}_{\frac{n}{2}}\right)(0:\tfrac{n}{2}-2)\\ 0&\underbrace{0\quad\ldots\quad 0}_{\frac{n}{2}-1}\end{pmatrix}}_{\text{Step 3}}.$ (8)
One can verify that this is indeed a particular three-step evaluation of Eq. (6) (left as an exercise for the interested reader). Indeed, Step 1 is the first level of the up-sweep part of Brent-Kung circuit, and Step 3 is the last level of the down-sweep part of the BK circuit. Step 2 is the recursive step. With a bit more work, we can show that the Ladner-Fischer circuits [9] can also be described as a specific evaluation of Kronecker decompositions.
The number of operations in Eq. (8) is well-known. Step 1 performs $\tfrac{n}{2}$ additions in parallel. Step 2 recursively computes a prefix sum for length $\tfrac{n}{2}$. Step 3 requires $\tfrac{n}{2}-1$ operations, giving a recursive formula for the size $S(n)=n-1+S(\tfrac{n}{2})$ operations. Unrolling the recursion, we get $S(n)=2n-\log(n)-2$ when $n$ is a power-of-two. The corresponding recursion  [^1] Indeed, Step 1 performs additions in parallel with depth one, the second step (recursion) has depth $D(n/2)$, and the last step performs additions in parallel. of the depth is $D(n)=1+D(n/2)+1$, which unrolls to $D(n)=2\log(n)-1$. In this scenario, the depth and sum give a suboptimal sum of $S(n)+D(n)=2n+\log(n)-3$, i.e., the resulting prefix circuit has $\log(n)+1$ deficiency.
Identifying the deficiency.
Note that in the update step of Eq. (7), we only need the first $\tfrac{n}{2}-1$ elements of $1_{2}^{\top}\bm{X}\bm{U}_{\frac{n}{2}}$. It turns out that this is precisely the “root” of the deficiency: every recursive call computes one element more than what is needed! With this observation, we can modify the recursion to only compute the first $\lceil\tfrac{n}{2}\rceil-1$ elements of $1_{2}^{\top}\bm{X}\bm{U}_{\frac{n}{2}}$:
  1. Step 1:
    $\bm{w}^{\top}\leftarrow\bm{1}_{2}^{\top}\bm{X}\in\mathbb{R}^{\lceil\frac{n}{2}\rceil}$ ($\lfloor\frac{n}{2}\rfloor$ binary additions).
  2. Step 2:
    $\bm{z}^{\top}\leftarrow\bm{w}^{\top}(0:n^{\prime})\bm{U}_{n^{\prime}}$, where $n^{\prime}:=\lceil\frac{n}{2}\rceil-1$.
  3. Step 3:
    Return:
    $\begin{pmatrix}\bm{X}(0,0:n^{\prime})&\bm{X}(0,\lceil\tfrac{n}{2}\rceil)\\ \bm{z}^{\top}(0:n^{\prime})&\bm{w}(\lceil\tfrac{n}{2}\rceil)\end{pmatrix}+\begin{pmatrix}0&\bm{z}^{\top}(0:n^{\prime}-1)&\bm{z}(n^{\prime})\\ 0&0\quad\ldots\quad 0&\bm{z}(n^{\prime})\end{pmatrix}$.
Evidently, this procedure does not only remediate the Brent–Kung deficiency, but it can also be generalized for any “block-size” $s\geq 2$, providing a remarkably simple, zero-deficiency prefix circuit family, with low depth.
A new zero-deficiency circuit family.
In Figure 2, we illustrate the proposed family of zero-deficiency prefix circuits. It consists of three layers:
  1. 1.
    The first layer partitions the $n$ inputs in blocks of size $s$, except the last block which has size $n\mod s$. A serial prefix circuit $L_{i}$ is used to compute the prefix of $\bm{x}(is:\min\{is+s,n\})$, for all $i=0,\ldots\lceil\tfrac{n}{s}\rceil-1$ simultaneously (in parallel).
  2. 2.
    The second layer takes as inputs the last outputs of every $L_{i}$, except the last one, and computes their prefix recursively. In this layer we can actually use any zero-deficiency circuit, and the final circuit will also have zero-deficiency. However, the depth, size, and fan-out, will be affected.
  3. 3.
    In the third layer, the outputs of second layer are used to finalize the remaining partial prefixes.
The analysis is summarized in Theorem 2. Notably, for $s=2$, the circuit has the same $2\log(n)+O(1)$ depth as Brent–Kung and other classic circuits, but, for every $n$, it achieves Snir’s lower bound S(n)=2n2D(n)S(n)=2n-2-D(n)!
Theorem 2.
For every pair of integers $n\geq 4$ and $s\in[2,n/2]$, there exists a LOGTIME-uniform family of recursive zero-deficiency prefix circuits with depth $sh_{s}^{*}(n)+r_{s}(n)-1\leq s\lceil\log_{s}(n)\rceil-2$ and fan-out $s$. If $n$ is a power of $s$, there is a single gate with fan-out $s+1$.
Proof.
We show by induction that a recursive construction achieves zero-deficiency. In the base case, if $\lceil\tfrac{n}{s}\rceil-1\leq s$, then we simply have a single serial prefix circuit with size n1n-1 and depth n1n-1, which has zero-deficiency. For the inductive step, we argue that if the claim holds for length $\lceil\tfrac{n}{s}\rceil-1$, then it also holds for length $n$. The first layer has size $n-\lceil\tfrac{n}{s}\rceil$ and depth s1s-1. The circuit of the second layer has zero-deficiency by the induction hypothesis, and therefore its depth and size sum to $2(\lceil\tfrac{n}{s}\rceil-1)-2=2\lceil\tfrac{n}{s}\rceil-4$. The final layer has depth one and size $n-s-\lceil\tfrac{n}{s}\rceil+2$. Adding everything together, the total depth and size sum to:
$\overbrace{n-\lceil\tfrac{n}{s}\rceil+s-1}^{\text{Layer 1}}+\overbrace{2\lceil\tfrac{n}{s}\rceil-4}^{\text{Layer 2}}+\overbrace{n-s-\lceil\tfrac{n}{s}\rceil+3}^{\text{Layer 3}}=2n-2,$
as desired. We can now upper bound the depth of the recursive circuit. We have that:
$D(n)=\begin{cases}s+D(\lceil\tfrac{n}{s}\rceil-1),&n>s,\\ n-1,&n\leq s.\end{cases}$
This quantity is always upper bounded by $s\lceil\log_{s}(n)\rceil-2$. To see this, note that $D(n)$ is equal to $sh^{*}_{s}(n)+r_{s}(n)-1$. Now, it always holds that $D(n)\leq D(s^{k})$, where $k=\arg\min\{j\ |\ s^{k}\geq n\}$. But if we replace $n$ with $s^{k}$ in the above, it holds that $h^{*}_{s}(s^{k})=k-1$ and $r_{s}(s^{k})=h_{s}^{(k-1)}=s-1$. Therefore, the total depth satisfies $D(s^{k})=s(k-1)+(s-1)-1=sk-2$. Since $k=\lceil\log_{s}(n)\rceil$, we finally obtain that $D(n)\leq D(s^{k})=s\lceil\log_{s}(n)\rceil-2,$ where the equality always holds when $n$ is a power of $s$.
To see that the construction can be verified in LOGTIME, note that given $n$, $s$, $i\in[n]$, $j\in[n]$, and $k\in[s\lceil\log_{s}(n)]$, as inputs, it is straightforward to answer in $O(\log(n))$ steps whether input $i$ is connected with input $j$ at level $k$. ∎
Figure 2: Zero-deficiency Kronecker prefix circuit.
Optimizing the block size.
For $s=2$, we obtain a zero-deficiency circuit with depth at most $2\lceil\log(n)\rceil-2\in 2\log(n)+O(1)$, which is the same as Brent–Kung for powers of two. However, $s=2$ does not always give the minimum achievable depth. Indeed, for larger values of $n$ we can sharpen the bound by setting $s=3$, which gives $D(n)\leq 3\lceil\log_{3}(n)\rceil-2\in\frac{3}{\log(3)}\log(n)+O(1)\approx 1.8928\log(n)+O(1)$. For any $n$, the actual minimum depth can be computed with dynamic programming: $\min D(n)=\min_{2\leq s\leq n/2}\{s+\min D(\lceil\frac{n}{s}\rceil-1)\}$.
IV Quantum adders
Here we mention how to apply the proposed prefix circuits to quantum adders. Such circuits have received significant attention the last three decades since they serve as building blocks for important problems such as discrete logarithm and integer factoring [27, 28, 29, 30, 31, 32, 4, 5]. Here we follow the methodology of [4, 5], where the main idea is to use a prefix circuit for carry propagation/generation (carry look-ahead framework). Circuits are evaluated based on the following metrics:
Toffoli count
The total number of Toffoli gates.
Toffoli depth
The maximum number of Toffoli gates in any path from an input to an output.
Qubit count
The total number of auxiliary qubits.
The goal is to minimize the Toffoli depth/count by carefully overlapping different independent Toffoli layers, while still using as few auxiliary qubits as possible. Figure 3 shows three cases where Toffoli gates are used in the quantum circuit for propagation and generation. We refer also to [5, Sections 3 and 4] for a broad overview of the techniques and details of the specific operations.
Figure 3: Toffoli gates for carry propagation/generation.
KroneckerQuantumAdder(n,s)(n,s):
[…]
  2. $\bm{p}_{0}(i):=\bm{b}(i)\leftarrow\bm{a}(i)\oplus\bm{b}(i),$ $i\in[0:1:n-1]$.
[…]
  6. $3:$ Uncompute $\bm{p}_{1}$.
  7. $4:$ Uncompute $\bm{p}_{0}$ and finalize sum.
[…]
  1. if $n\leq s:$ /* Final Layer */
    1. for $i=1,\ldots,n-1:$
      1. $i:\bm{g}(i)\leftarrow\mathcal{T}\Big(\bm{g}(i-1),\bm{p}_{t-1}(i),\bm{g}(i)\Big)$
  2. else:
    1. Set $n^{\prime}=\left\lceil\tfrac{n}{s}\right\rceil-1$ and $I=\left[s-1:s:n^{\prime}-1\right]$.
[…]
  1. for $k=1,\ldots,s-2:$
    1. for $i\in[k+s:s:n-1]:$
      1. $k:\bm{g}(i)\leftarrow\mathcal{T}\Big(\bm{g}(i-1),\bm{p}_{t-1}(i),\bm{g}(i)\Big)$
[…]
  2. Set k=s1k=s-1
  3. $k:\bm{g}(i)\leftarrow\mathcal{T}\Big(\bm{g}(i-1),\bm{p}_{t-1}(i),\bm{g}(i)\Big)$ $i\in[s-1:s:n-1]$
[…]
  1. for $i\in[s:s:n-2]:$
    1. for $k\in\big[i:\min\{i+s-2,n-2\}\big]:$
      1. :*: $\bm{z}(is+k)\leftarrow\bm{g}(i-1)\oplus\bm{z}(is+k)$
      2. $1:\bm{g}(k)\leftarrow\mathcal{T}\Big(\bm{z}(is+k),\bm{p}_{t}(k),\bm{g}(k)\Big)$.
  2. $1:\bm{g}(n-1)\leftarrow\mathcal{T}\Big(\bm{g}(s\lfloor\tfrac{n}{s}\rfloor),\bm{p}_{t}(n-1),\bm{g}(n-1)\Big)$.
  3. $1:$ Uncompute $\bm{p}_{t+1}$ and $\bm{z}$.
Figure 4: Algorithm to construct quantum adder. The lines that are numbered indicate the corresponding Toffoli layer (layers marked as :*: do not contain Toffoli gates).
Table 3: Comparison of known quantum adders with respect to Toffoli depth, count, and number of auxiliary qubits. BK:=Brent–Kung, Other $\in$ {Sklansky, Kogge–Stone, Han–Carlson, Ladner–Fischer}.
Quantum Adder Toffoli Count Toffoli Depth Qubit Count
BK [4, 5] $O(n)$ $2\log(n)+O(1)$ $O(n)$
Other, Strategy 1 [5] $O(n\log(n))$ $2\log(n)+O(1)$ $O(n\log(n))$
Other, Strategy 2 [5] $O(n\log(n))$ $\log(n)+O(1)$ $O(n\log(n))$
Thm. 3, $s=3$ $O(n)$ $1.893\log(n)+O(1)$ $O(n)$
[…]
Theorem 3.
[…]
Proof.
We start bounding the qubit count. The algorithm uses nsn-s auxiliary qubits for the register $\bm{z}$ and at most $O(n/s^{t})$ qubits for each $\bm{p}_{t}$, for $t=1,2,\ldots,\lceil\log_{s}(n)\rceil+1$. This gives a total of $O(n)$ auxiliary qubits.
For the Toffoli depth, there are two layers of Toffoli gates before the first recursive call, and two layers after (for uncomputation). $\textbf{Recurse}(\cdot)$ has Toffoli depth at most $s\lceil\log_{s}(n)\rceil-2$, which is the depth of the Kronecker prefix circuit. Therefore, the total Toffoli depth is at most $4+s\lceil\log_{s}(n)\rceil-2=s\lceil\log_{s}(n)\rceil+2$.
The total number of Toffoli gates is bounded as follows. In the first two propagation layers, $\bm{p}_{0}$ and $\bm{p}_{1}$, there are a total of $3n/2$ gates, and additional $n/2$ gates for the un-computation of $\bm{p}_{1}$. In each recursive step $t=1,\ldots,\lceil\log_{s}(n)\rceil-1$, there are at most $n/s^{t}$ gates for $\bm{g}^{(t+1)}$, at most $n/s^{t+1}$ gates for $\bm{p}^{(t+2)}$, at most $n/2^{t}$ gates for $\bm{g}^{(t)}$, and finally at most $n/2^{t+1}$ gates for the uncomputation of the $\bm{p}^{(t+2)}$. This gives a total of at most $3n/2^{t}$ gates at level $t$. Summing for all $t=1,\ldots,\lceil\log(n)\rceil-1$, we obtain at most $3n$ Toffoli gates. ∎
V Conclusion
In this work we revisited algorithms and circuits for the prefix sum problem, through the lens of linear algebra. By decomposing triangular all-ones matrices as the sum of two Kronecker products, we were able to describe a new family of recursive zero-deficiency prefix circuits. These circuits are parametrized by an integer $s$ (the block-size). By choosing $s$ appropriately, we can obtain circuits with reduced depth and/or fan-out compared to existing zero-deficiency families. As an application, these techniques were used to construct quantum adders with reduced Toffoli depth and/or size, compared to existing ones.